### v2

# Butterfly in Spacetime: Inherent Instabilities in Stable Black Holes

Abstract
This paper numerically studies if the stability of a stable black hole is robustness against the small perturbation on geometry near its event horizon. As a toy model, it encodes the such perturbation into deformations of Regge-Wheeler potential. It considers three different types of local deformations-the negative static bump potential, the stochastic potential and bump potential modulated by time function in low frequency limit. Our numerical results show that infinitesimal local deformations on Regge-Wheeler potential near the horizon can overturn stability of an arbitrarily strongly stable black hole, implying that late-time behavior of a stable black hole is extremely sensitive to geometry near horizon. Specially, certain deformations that stabilize systems in flat backgrounds can destabilize otherwise stable black holes. It also shows that horizon-induced redshift transforms near-horizon quantum fluctuations into classical-scale stochastic deformations capable of triggering instability, implying that even an isolated black hole cannot keep stable in extended timescales.
I Introduction
In the field of black hole (BH) physics, the analysis of BH stability stands as a fundamental and critical subject, as it reveals whether BHs can maintain their existence in the universe over extended periods. The quasi-normal modes (QNMs), meanwhile, serve as one of the essential physical quantities for exploring and understanding the stability characteristics of BHsChandrasekhar and Detweiler (1975); Nollert and Price (1999); Kokkotas and Schmidt (1999); Berti et al. (2009); Konoplya and Zhidenko (2011). Unlike normal modes in close system, the QNMs are a series of discrete complex modes. The sign of the imaginary parts indicates whether the spacetime is stable or not. In general, QNMs indicating unstable spacetime describe a system where phase transition happens, such as spontaneous scalarization of compact objectSilva et al. (2018); Herdeiro et al. (2018); Doneva et al. (2024); Xu et al. (2024); Xiong and Li (2024), while QNMs describing stable spacetime play an important role in ringdown signal of gravitational waves in the last period of two astrophysical compact object mergeCreighton (1999); Dreyer et al. (2004); Berti et al. (2006); Gossan et al. (2012); Berti et al. (2018); Wang et al. (2024).
Recently, an interesting feature of QNMs, even the fundamental modes, has been reported that they are sensitive to small alterations of background, which called spectrum instabilityNollert (1996); Cheung et al. (2022); Cardoso et al. (2024); Qian et al. (2025). For example, when the effective potential is altered by adding a small positive “bump” away from the location of photon ring, the quasi-normal modes (QNMs) of a Schwarzschild BH are completely alteredCheung et al. (2022). The sensitivity of QNMs seems to imply that the QNMs are not appropriate observables in ringdown stage. However, the ringdown signal appear to be robust by the spectral instabilityDaghigh et al. (2020); Berti et al. (2022); Yang et al. (2024); Ianniccari et al. (2024); Rosato et al. (2024). During the early ringdown stage, the signal in altered background exhibits no substantial divergence from the original signal. In the late ringdown stage, while a discernible discrepancy arises between the two, the magnitude of this difference is directly proportional to the intensity of the bump and will exponentially decays to zero. Notably, both signals undergo rapid decay to zero, maintaining consistency in their asymptotic behavior. The ringdown signal is nonsensitive against the small alterations of effective potential in these literatures.
When addressing the topic of BH stability, the focus is usually on whether a probe field exhibits exponential growth in a fixed spacetime background. Numerous studies have demonstrated that Schwarzschild BH remains stable under perturbations from scalar fields, Dirac fields, and spin-2 fieldsRegge and Wheeler (1957); Chandrasekhar (1998). Due to the complex environment and unavoidable classical/quantum fluctuations of real BHs, no models in these studies can fully characterize any real BH in the universe. However, these investigations are still regarded as highly valuable. A key reason is the prevailing belief that the conclusion about BH stability is robust against small alterations of the spacetime metric-that is, sufficiently small alteration on the metric would not disrupt the stability of an initially stable BH. In other words, if a theoretical BH described by metric $g_{\mu\nu}$ is stable, then we believe the real BH described by metric $\tilde{g}_{\mu\nu}$ will also be stable if the difference between $\tilde{g}_{\mu\nu}$ and $g_{\mu\nu}$ is “small” enough. However, our results in this work suggest that this “taken-for-granted” assumption may not be valid. Using the Schwarzschild BH as a concrete example, we consider if its stability is sensitive to the modification of geometry near horizon. As a toy model, we encode such modification into the deformation of Regge-Wheeler potential. We analyze the long-time behavior of probe scalar field when a small negative or zero-mean random local effective potential is added sufficiently close to the horizon, extending the analysis beyond the positive bump scenario investigated inCheung et al. (2022); Destounis and Duque (2024). In the time domain, we numerically show that the spacetime becomes unstable when the deformation is sufficiently close to the horizon. Particularly, certain perturbed potentials that stabilize systems in flat backgrounds can destabilize otherwise stable black holes. We conclude that, at least under static spherical symmetry, the stability of a BH does not exhibit robustness. Two nearly identical BHs may exhibit completely different stability properties at late times. It will also show that horizon-induced redshift transforms near-horizon quantum fluctuations into classical-scale stochastic local deformations on the metric, which can trigger instability of the BH and cause the BH gradually to lose energy over time. In this paper, we use natural units $\hbar=c=G=1$.
II Effective potential and its deformations
We take the original spacetime to be Schwarzschild BH, of which the line element is
$\text{d}s^{2}=-h(r)\text{d}t^{2}+\frac{\text{d}r^{2}}{h(r)}+r^{2}\text{d}\Omega^{2},~{}~{}~{}h(r)=1-r_{h}/r\,,$ (1)
where $r_{h}$ is the radius of event horizon, $\text{d}\Omega^{2}$ stands for line element of unit sphere. In time domain, we consider equation of motion with respect to the s-wave scalar perturbation $\psi(t,r_{*})$ in the Schwarzschild spacetime
$\frac{\partial^{2}\psi}{\partial t^{2}}-\frac{\partial^{2}\psi}{\partial r_{*}^{2}}+V_{\mathrm{eff}}\psi=0\,.$ (2)
The tortoise coordinate $r_{*}$ is defined as $\text{d}r_{*}/\text{d}r=1/h$, implying that $r\to r_{h},r_{*}\to-\infty$ and $r\to\infty,r_{*}\to\infty$. The effective potential in Eq. (2) is just the Regge-Wheeler potential $V_{\mathrm{eff}}=V_{\mathrm{RW}}=hh^{\prime}/r$. The $V_{\mathrm{RW}}$ is nonnegative, so the Schwarzschild BH is stable against the scalar perturbation, which has been confirmed inRegge and Wheeler (1957).
We now add a deformation term into the effective potential to capture possible environmental effects presenting around the BH system. Detailed discussion on how the environments change Eq. (2) will much involved and we here simply assume that they are all encoded into a deformation on effective potential
$V_{\mathrm{eff}}=V_{\mathrm{RW}}(r_{*})+\epsilon V_{p}(r_{*}-a,t)\,,$ (3)
where $V_{p}(x,t)$ is a function that localizes around at $x=0$ and decays in spatial direction at least as fast as $V_{\mathrm{RW}}$. Note that $V_{p}(x,t)$ may not decay in time. Though it might not characterize details of any realistic setup and the dynamical evolution of matter and gravity might also need to be considered, it is usually used as a toy model and a beginning point in understanding how environments change the stability of the BH Cheung et al. (2022); Barausse et al. (2014); Chung et al. (2021); Capuano et al. (2024).
In following discussions, we will specify $r_{h}=1$. We present our numerical results for solving Eq. (2) in time domain with initial data corresponding to Gaussian wave packet
$\psi(0,r_{*})=\exp\left(\frac{r_{*}^{2}}{20}\right)\,,\quad\frac{\partial\psi}{\partial t}(0,r_{*})=0\,.$ (4)
For spatial grid $r_{*}$, we adopt forth-order central difference method for spatial derivative and integrate the Eq. (2) using forth-order Rugge-Kutta method for $t$. Here we will consider three different types of potential $V_{p}(r_{*}-a,t)$.
III Static negative bump
The first one is the static “bump” potential $V_{p}(x,t)=V_{p}(x)$, where $V_{p}(x)$ has certain sign and compact support. When $V_{p}(x)$ is positive, it has been well known that such deformation will lead to “spectrum instability”. However, this kind of “instability” is not the instability of BH itself, i.e., the system will still settle down into a static BH without scalar hair. Since the bumps are added by hand, there is not proper justification for a positive or negative sign. Particularly, near the event horizon, the quantum effects of gravity and matter may lead to exotic matter and may raise negative “bump”. The negative “bump” can even appear if we impose dominate energy condition (see one example in appendix). The negative "bump" could resemble a tachyonic field Cardoso et al. (2009); Baibhav et al. (2023) or be inherent in various modified gravity theories, such as the Einstein-Gauss-Bonnet theory Cao et al. (2024). If the bump is close to a stable BH, the situation become complicated. We expect a competition with the absorption at the horizon. For strong enough negative bump, this feature might give rise to spacetime instabilitiesCardoso et al. (2009); Baibhav et al. (2023); Cao et al. (2024). Naively, one may expect that the BH will be still stable if the original BH is strongly enough but the negative “bump” is small enough. Surprisingly, our numerical results show that, if the infinitesimal negative “bump” is close to event horizon enough, it can turnover arbitrarily stable BH.
We specify the negative bump to be Gaussian form around center $r_{*}=a$
$V_{p}(r_{*}-a)=-\exp\left[-\frac{(r_{*}-a)^{2}}{16}\right]\,.$ (5)
We numerically investigate how this negative bump change the long-time behavior of scalar perturbation when it is placed close to even horizon.
Figure 1: Left panel: Time evolution of $\psi$ at $x=a$ for different values of $\epsilon$ and $a$. Here $t_{0}$ is defined as the first time when $\psi(a,t)$ is at its local maximum value. Right panel: Relationship between $\epsilon$ and $|a_{c}|$. The error bars are determined by the bounds of bisection. The red line is given by equation $\ln|a_{c}|=-\ln|\overline{V}_{p}|-\ln\epsilon$ (not by fitting).
The left panel of Fig. 1. shows the results as we move perturbed potential towards event horizon by choosing $\epsilon=10^{-2}$ and $4\times 10^{-3}$. It shows that for small value of $a$ where the deformations are confined near the photon ring, $\psi$ will decay to approach zero in the last stage, confirming that the BH is stable. However, for the case $\epsilon=10^{-2}$ we found that there exist a critical distance $a_{c}\approx-15$ of deformation to trigger the BH instability. When $a<a_{c}$, the perturbation field $\psi$ will experience exponential growth at late time. This exponential amplification signals an onset of instability, indicating that the perturbation field will drive the BH towards an unstable state during the later stage of its evolution.
If we decrease the amplitude $\epsilon$, then we find that the critical distance will also move towards horizon. Using the method of bisection, we can approximately find the relationship between $\epsilon$ and $a_{c}$, which is shown in the right panel of Fig. 1. Particularly, we find following asymptotically relationship between $\epsilon$ and $a_{c}$
$a_{c}=1/(\overline{V}_{p}\epsilon)\,$ (6)
when $\epsilon$ is small enough. Here $\overline{V}_{p}=\int_{-\infty}^{\infty}V_{p}(x)\text{d}x$. We have tested various differen types of negative bumps and find that Eq. (6) is true in all cases. Our numerical results imply that the instability will always be triggered if the negative bump is sufficiently close to event horizon no matter how small the amplitude is.
IV stochastic local potential
In the second form we suppose $V_{p}(x,t)$ to be local stochastic function with zero mean. This in some sense is closer to real astrophysical environments compared with a fixed-sign perturbation. Many astrophysical BHs are surrounded by dynamical evolutional matter. The matter supplies non-static stochastic deformations on the geometry of BH. As a toy model, the effects of random matter then could be mimicked by random non-static modification on the Regge-Wheeler potential. More importantly, this kind of modification seemingly is inevitably even for an isolated BH without any matter surrounding it, since it can come from the local quantum fluctuations. Based on equivalent principle, the wavelength $\lambda_{p}$ and period $t_{p}$ of local quantum fluctuation measured by local inertial observers are Planck scales $\lambda_{p}=t_{p}\sim\mathcal{O}(l_{\rm P})$. These quantum fluctuations will lead that the geometry of spacetime varies in the local proper time scale $t_{p}$ and local proper distance scale $\lambda_{p}$. Usually, such extreme ultraviolet variaions of metric have no dramatic effect on classical physics. But the appearance of horizon makes thing different due to the strong redshift.
To illustrate that, we consider the quantum fluctuation happening in the near-event-horizon region $r=r_{h}+\delta r$ (measured in coordinate gauge (1)). The line element in terms of tortoise coordinate in this region then will be $\text{d}s^{2}=h(r)(-\text{d}t^{2}+\text{d}r_{*}^{2})\approx h^{\prime}(r_{h})\delta r(-\text{d}t^{2}+\text{d}r_{*}^{2})$ such that one finds $\lambda_{p}=\sqrt{h^{\prime}(r_{h})\delta r}\lambda=\sqrt{h^{\prime}(r_{h})\delta r}T$. Here $\lambda$ and $T$ denote the wavelength and period in tortoise coordinate. The picture of classical horizon would fail if one is close to it near Planck scale. Thus, the proper distance between $r=r_{h}+\delta r$ and $r=r_{h}$ can at most as same as Planck scale. This means $\delta r/\sqrt{h(r)}\approx\delta r/\sqrt{h^{\prime}(r_{h})\delta r}\sim\mathcal{O}(\lambda_{p})$. One thus find that $T=\lambda\sim 1/h^{\prime}(r_{h})\sim\mathcal{O}(r_{h})$. Therefore, from the perspective of tortoise coordinate, near horizon local quantum fluctuations would induce metric or Regge-Wheeler potential to varies in infrared scale. It needs to emphasize again that, in the case of absent horizon or in the region far way from horizon, the local quantum fluctuations will have no essential contributions on classical physics in infrared scale.
In order to characterize the randomness of variation on Regge-Wheeler potential, we specify
$V_{p}(x,t)=W(x)U(t)\,,$ (7)
where $W(x)$ is random function localizing around $x=0$ and $U(t)$ is a temporal random function. The function $W$ and $U$ satisfy
$\overline{W}=\int^{\infty}_{-\infty}W(x)\text{d}x=0,~{}\overline{U}=\lim_{t_{m}\rightarrow\infty}\frac{1}{t_{m}}\int^{t_{m}}_{-t_{m}}U(t)\text{d}t=0\,.$ (8)
The differences between $W$ and $U$ in Eq. (8) is because that $W$ is required to localize around $x=0$ but function $U$ is assumed to have nonzero distribution in whole $t\in\mathbb{R}$. In addition, the spectrum (after of Fourier transformations) of $W$ and $U$ should mainly distribute in the region that are not larger than the order $\mathcal{O}(1/r_{h})$.
Figure 2: Left panel: Time evolution of $\psi$ at $x=a$ for different values of $\epsilon$ and $a$. Right panel: Relationship between $\epsilon$ and $|a_{c}|$. The red line is $\ln|a_{c}|=-2\ln\epsilon+0.02$, which is obtained by fitting the most left 7 points.
As a concrete example, we choose pseudo-random functions without any explicit periodicity and symmetries such that $U(t)=\sin(t/2)+\cos(t/\sqrt{2})-\sin(t/\pi)$ and
$W(x)=w_{0}\left[1-\tanh(4(x-2))\tanh(4(x+2))\right]Q(x)\,.$ (9)
with
$Q(x)=\sin x+\cos\sqrt{2}x+\cos\pi x-\sin\pi x+c_{0}\,.$ (10)
Here $c_{0}$ and $w_{0}$ are chosen so that $W$ has zero average and normalization condition $\max|W|=1$ is satisfied. The detailed forms of Eq. (9) is not important. We choose functions (9) and (10) just as a concrete example. The results are shown in Fig. 2. We still find that there is a critical $a_{c}$ for every given amplitude $\epsilon$. The instability will be triggered if $a<a_{c}$. The method of bisection gives asymptotical behavior
$a_{c}\propto 1/\epsilon^{2}$ (11)
for sufficiently small $\epsilon$. We have tested many different functions of $W$ and $U$ and find that Eq. (11) is universal. The above numerical results still imply that the instability will be triggered if the random nonstatic deformation on the effective potential is sufficiently close to event horizon no matter how small the amplitude is.
V Low-frequency instability
In above two examples, perturbed potentials $V_{p}$ themselves will trigger the instability if we turn off the Regge-Wheeler potential $V_{\text{RW}}$. Thus, roughly speaking, the instability can be understood as the competition between the destabilizing caused by perturbed potential $V_{p}$ and the stabilizing caused by BH. The following example will show that this is not the whole story. We specify the perturbed potential $V_{p}$ to be
$V_{p}(r_{*}-a,t)=\exp\left[-\frac{(r_{*}-a)^{2}}{\sigma^{2}}\right][\sin(wt)+\sin(\pi wt)]\,,$ (12)
i.e, we consider a time-dependent perturbed potential that has Gaussian distribution in space but pseudo-random in time. We here focus on the situation $w\ll\mathcal{O}(1/r_{h})$.
Figure 3: Left panel: Time evolution of $\psi$ at $x=a$ in different situations with $w=0.02$ and $\sigma=4$. Right panel: Relationship between $\epsilon$ and $|a_{c}|$ for $w=0.001$ and $\sigma=4$. The red line is $\ln|a_{c}|=-\ln\epsilon-0.93$, which is obtained by fitting the most left 6 points.
As one concrete example, we take $\sigma=4$ and $w=0.02$. The results are shown in Fig. 3. If we turn off the BH potential (i.e. setting $V_{\text{RW}}=0$) but keep the perturbed potential $V_{p}$ by setting $\epsilon=0.02$, we find that the scalar probe field approximately exponentially decays into zero. Thus, the perturbed potential $V_{p}$ itself will stabilize the system. When we combine perturbed potential and Regge-Wheeler potential together, we find that there is critical distance $a_{c}\approx-35$ and the instability will be triggered if $a<a_{c}$. In the region $w\in[0.001,0.02]$ the results are similar. The numerical accuracy restricts us to explore smaller $w$. For sufficiently small $w$, saying $w=0.001$ for example, the method of bisection gives us following asymptotical behavior
$a_{c}\propto 1/\epsilon$ (13)
for sufficiently small $\epsilon$. Our numerical results suggest that, for the case $w\rightarrow 0^{+}$ the infinitesimal $\epsilon$ can still trigger the instability if the perturbed potential is sufficiently closed to the horizon. This effect appears unique to black holes as these same potentials do not destabilize Minkowski spacetime.
VI conclusion
Astrophysical BHs are inevitably surrounded by matter distributions, and even isolated BHs experience quantum fluctuations. These classical and quantum perturbations induce stochastic deformations in spacetime geometry. This study addresses a critical question: whether nominally stable BHs retain their stability when the geometry are deformed by such small perturbations. To model this effect, we map the deformations onto modifications of the effective potential governing s-wave scalar fields. Numerical analysis provides compelling evidence that BH stability can be disrupted by infinitesimal negative or stochastic modifications localized near the event horizon. Notably, certain perturbed potentials that stabilize systems in flat backgrounds can destabilize otherwise stable BHs. While our focus is on the event horizon, analogous behavior may occur near cosmological horizons in asymptotically de Sitter BHs. Our findings suggest that the stability of a stable BH is not robust against the tiny deformation of near-horizon geometry. Infinitesimal deformation of near-horizon geometry can turnover the stability of a stable BH. Due to the inevitable fluctuations of near horizon geometry, this intrinsic instability fundamentally constrains our ability to predict BH dynamics over extended timescales.
As a toy model and very preliminary research, we here only considered static spherically symmetric case and numerically simulated the system in time domain. It would be interesting to consider Kerr black hole and study how spin would affect the phenomenon discovered here. Particularly, is there any relationship to turbulence of black holes discovered in Ref. Yang et al. (2015)? Could we use frequency analysis to verify the results and supply deeper understandings on the underlying mechanism?  [^1] The analysis of frequency domain for static negative bump and static stochastic potential will be addressed in an upcoming publication. Why does bump potential modulated by a time function in low frequency limit stabilize the flat spacetime but destabilize the black hole? More realistically, could we quantitatively estimate the disturbance intensity $\epsilon$ generated by the actual environment (such as accretion disks, dark matter halos) or quantum fluctuations, and whether the position $a$ satisfies the condition of $a<a_{c}$? We hope these questions could be addressed in the future through numerical simulations and further theoretical works.
At the classical level, we emphasize that our results do not imply instability of the vacuum Schwarzschild solution, as no source exists to generate stochastic deformations on metric. For a BH surrounded by matter, our results implies that the system cannot keep static configuration for a long time in astrophysical settings, since uncontrollable stochastic matter fluctuations near the horizon will lead stochastic deformations on the geometry and inevitably trigger the instability identified here. The timescale that this instability becomes considerable depends on amplitude of the deformation. The probe scalar field studied here can act as a conduit, transferring energy from the environment to the BH and spatial infinity. Consequently, the final state evolves toward a Schwarzschild BH in vacuum, with increased horizon area and energy. However, if we posit an underlying quantum-gravitational framework beneath classical general relativity, our analysis indicates that even vacuum Schwarzschild BHs become unstable. This suggests that Schwarzschild BHs in a quantum-gravitational framework would gradually lose energy over time, potentially offering a new perspective on BH evaporation. Further investigation is needed to establish a direct connection and we hope some useful advances will be obtained in the future.
Acknowledgements.
We thank Prof. Vitor Cardoso for sharing considerations on relationship between “negative bump” and instability of spacetime. We thank Yiqiu Yang for helpful discussions. Z. F. Mai is supported by National Key R&D Program of China (2024YFA1611700). R. Q. Yang is supported by the National Natural Science Foundation of China under Grant No. 12375051. This work is also supported by the Guangxi Talent Program (“Highland of Innovation Talents”).
Appendix A Negative bump potential and dominate energy condition
In this appendix, we will give a concrete example that Regge-Wheeler potential has a local negative bump but dominate energy condition is satisfied.
Let us first begin from general static spherical asymptotically flat BH in tortoise coordinates, of which the metric reads
$\text{d}s^{2}=h(x)(-\text{d}t^{2}+\text{d}x^{2})+r(x)^{2}\text{d}\Omega^{2}\,.$ (14)
We here assume that the event horizon locates at $x=-\infty$, where we have $r(-\infty)=r_{h}$ and $h(-\infty)=0$. The function $h(x)$ has asymptotically expression $h(x)=1-2M/x+\cdots$ with the total mass $M$ of the spacetime and $r^{\prime}(\infty)=1$. The Regge-Wheeler potential $V_{\mathrm{RW}}$ for s-wave scalar field then reads
$V_{\mathrm{RW}}=r^{\prime\prime}/r\,.$ (15)
In static spherically symmetric case, the energy momentum tensor ${T^{\mu}}_{\nu}$ has a form ${T^{\mu}}_{\nu}=\text{diag}[-\rho(x),p_{x}(x),p_{T}(x),p_{T}(x)]$. The dominate energy condition then requires
$\rho\geq|p_{x}|,~{}\rho\geq|p_{T}|\,.$ (16)
The Einstein’s equation gives following independent equations.
$h^{\prime}$ $=$ $\frac{1+8\pi r^{2}p_{x}}{rr^{\prime}}h^{2}-\frac{hr^{\prime}}{r}\,,$ (17)
$rr^{\prime\prime}$ $=$ $h-r^{\prime 2}-4\pi r^{2}(\rho-p_{x})h\,\,.$ (18)
We combine Eq. (18) and (15) to obtain
$V_{\mathrm{RW}}=\frac{h-r^{\prime 2}-4\pi r^{2}(\rho-p_{x})h}{r^{2}}\,.$ (19)
For Schwarzschild BH, we have $\rho=p_{x}=0,h=1-2M/r$ and $r^{\prime}=h$. We then recover the result $V_{\mathrm{RW}}=2M(1-2M/r)/r^{3}$.
We now consider the near horizon region in general case. From Eqs. (17) and (18) one can find that there is no constraint on the values of $\rho-p_{x}$ and $p_{x}$. The values of $\rho$ and $p_{x}$ could be considerably large near horizon. Particularly, if an astrophysical BH is surrounded by sufficient matter localization around $x\sim a\ll-r_{h}$ and the dominate energy condition is satisfied, then it is possible
$4\pi r^{2}(\rho-p_{x})>1-r^{\prime 2}/h\,$ (20)
and we then see that the sign of $V_{\mathrm{RW}}$ would be negative.
As a concrete example to confirm above statement, we take $p_{x}(x)=-\rho(x)$ and
$\rho(x)=\left\{\begin{split}b>0,~{}~{}~{}|x-a|<1/8\\ 0,~{}~{}~{}|x-a|>1/8\end{split}\right.\,$ (21)
as an example. To solve the Einstein equation, we note that the system is in vacuum and the solution is a Schwarzschild BH with Schwarzschild radius $r_{h}$ in the region $x<a-1/8$, of which the metric reads
$x=r_{h}+r_{h}\ln(1-r_{h}/r),~{}~{}r^{\prime}=h=1-r_{h}/r\,.$ (22)
We then integrate equations (17) and (18) into infinity with the initial values at $x=a-1/8$ given according to Eq. (22).
Figure 4: Regge-Wheeler potential for $a=-20,r_{h}=1$ and $b=0.2$. Here $x_{m}\approx-11.09$ is the position where $V_{\mathrm{RW}}$ reaches its maximum.
In Fig. 4 we give a numerical result when we choose $a=-20,r_{h}=1$ and $b=0.2$. We see that there is negative bump around $x=a$. Thus, the negative bump potential can appear even if dominate energy condition is satisfied. We expect that, in general, localized spherical matter shells will introduce similar bumps in the Regge-Wheeler potential, possibly causing instabilities discussed in our main text.

### v3

@@
 Abstract
-This paper numerically studies if the stability of a stable black hole is robustness against the small perturbation on geometry near its event horizon.
+This paper numerically studies if the stability of a stable black hole is robust against the small perturbation on geometry near its event horizon.
 As a toy model, it encodes the such perturbation into deformations of Regge-Wheeler potential.
 It considers three different types of local deformations-the negative static bump potential, the stochastic potential and bump potential modulated by time function in low frequency limit.
-Our numerical results show that infinitesimal local deformations on Regge-Wheeler potential near the horizon can overturn stability of an arbitrarily strongly stable black hole, implying that late-time behavior of a stable black hole is extremely sensitive to geometry near horizon.
+Our numerical results show that infinitesimal local deformations on Regge-Wheeler potential near the horizon can overturn stability of a stable black hole, implying that late-time behavior of a stable black hole is extremely sensitive to geometry near horizon.
 Specially, certain deformations that stabilize systems in flat backgrounds can destabilize otherwise stable black holes.
@@
 In the field of black hole (BH) physics, the analysis of BH stability stands as a fundamental and critical subject, as it reveals whether BHs can maintain their existence in the universe over extended periods.
+The stability of a BH addresses the question: if it is subjected to a small perturbation, will the perturbations decay with time, allowing the BH to return to its original state, or will they grow uncontrollably and destroy the configuration altogether?
+The stability problem has deep implications for the mathematical structure of general relativity, cosmic censorship, and the no-hair conjecture.Chandrasekhar (1983).
 The quasi-normal modes (QNMs), meanwhile, serve as one of the essential physical quantities for exploring and understanding the stability characteristics of BHsChandrasekhar and Detweiler (1975); Nollert and Price (1999); Kokkotas and Schmidt (1999); Berti et al.
 (2009); Konoplya and Zhidenko (2011).
-Unlike normal modes in close system, the QNMs are a series of discrete complex modes.
+Unlike normal modes in closed system, the QNMs are a series of discrete complex modes.
 The sign of the imaginary parts indicates whether the spacetime is stable or not.
@@
 (2024); Xu et al.
-(2024); Xiong and Li (2024), while QNMs describing stable spacetime play an important role in ringdown signal of gravitational waves in the last period of two astrophysical compact object mergeCreighton (1999); Dreyer et al.
+(2024); Xiong and Li (2024), while QNMs describing stable spacetime play an important role in ringdown signal of gravitational waves in the last period of two astrophysical compact object mergerCreighton (1999); Dreyer et al.
 (2004); Berti et al.
@@
 (2024).
-Recently, an interesting feature of QNMs, even the fundamental modes, has been reported that they are sensitive to small alterations of background, which called spectrum instabilityNollert (1996); Cheung et al.
+Recently, an interesting feature of QNMs, even the fundamental modes, has been reported that they are sensitive to small alterations of background, which is called spectrum instabilityNollert (1996); Cheung et al.
 (2022); Cardoso et al.
@@
 For example, when the effective potential is altered by adding a small positive “bump” away from the location of photon ring, the quasi-normal modes (QNMs) of a Schwarzschild BH are completely alteredCheung et al.
-(2022).
-The sensitivity of QNMs seems to imply that the QNMs are not appropriate observables in ringdown stage.
-However, the ringdown signal appear to be robust by the spectral instabilityDaghigh et al.
+(2022).In addition, the spiral formula of spectrum instability can be semi-analytically derived in double rectangular barriers using perturbation theory (See the Supplemental material in Motohashi (2025)) The sensitivity of QNMs seems to imply that the QNMs are not appropriate observables in ringdown stage.
+However, the ringdown signal appears to be robust by the spectral instability Daghigh et al.
+(2020); Kyutoku et al.
 (2023); Berti et al.
@@
 (2024).
+Specifically, it was pointed out that the spectrum instability can hardly affect the greybody factor and on the ringdown signal.
+Further, the phase shift of the greybody factor could carry more physical information when studying the environmental effects on QNMs and the ringdown signal Kyutoku et al.
+(2023).
 During the early ringdown stage, the signal in altered background exhibits no substantial divergence from the original signal.
@@
 Notably, both signals undergo rapid decay to zero, maintaining consistency in their asymptotic behavior.
-The ringdown signal is nonsensitive against the small alterations of effective potential in these literatures.
-When addressing the topic of BH stability, the focus is usually on whether a probe field exhibits exponential growth in a fixed spacetime background.
+The ringdown signal is nonsensitive against the small positive alterations of effective potential.
+From a more macroscopic perspective, the stability of BHs remains unchallenged by this minor alteration on effective potential.
+When addressing the topic of BH stability, the focus is linear stability of BHs, which usually focuses on whether a probe field exhibits exponential growth in a fixed spacetime background.
+Linear stability analysis of BHs involves solving wave-like equations for perturbations of the metric or matter fields Chandrasekhar (1983).
 Numerous studies have demonstrated that Schwarzschild BH remains stable under perturbations from scalar fields, Dirac fields, and spin-2 fieldsRegge and Wheeler (1957); Chandrasekhar (1998).
@@
 In other words, if a theoretical BH described by metric $g_{\mu\nu}$ is stable, then we believe the real BH described by metric $\tilde{g}_{\mu\nu}$ will also be stable if the difference between $\tilde{g}_{\mu\nu}$ and $g_{\mu\nu}$ is “small” enough.
+This naive idea is seemingly supported by some models in studying “spectrum instability” Destounis and Duque (2024), since the perturbed potential in those models do not change instability of BHs in time domain.
 However, our results in this work suggest that this “taken-for-granted” assumption may not be valid.
 Using the Schwarzschild BH as a concrete example, we consider if its stability is sensitive to the modification of geometry near horizon.
-As a toy model, we encode such modification into the deformation of Regge-Wheeler potential.
+As a toy model, we encode such modification into the alteration of Regge-Wheeler potential.
 We analyze the long-time behavior of probe scalar field when a small negative or zero-mean random local effective potential is added sufficiently close to the horizon, extending the analysis beyond the positive bump scenario investigated inCheung et al.
-(2022); Destounis and Duque (2024).
-In the time domain, we numerically show that the spacetime becomes unstable when the deformation is sufficiently close to the horizon.
-Particularly, certain perturbed potentials that stabilize systems in flat backgrounds can destabilize otherwise stable black holes.
-We conclude that, at least under static spherical symmetry, the stability of a BH does not exhibit robustness.
-Two nearly identical BHs may exhibit completely different stability properties at late times.
-It will also show that horizon-induced redshift transforms near-horizon quantum fluctuations into classical-scale stochastic local deformations on the metric, which can trigger instability of the BH and cause the BH gradually to lose energy over time.
+(2022).
+In the time domain, we numerically show that the spacetime becomes unstable when the alteration is sufficiently close to the horizon.
+Particularly, certain perturbed potentials that stabilize systems in flat backgrounds can destabilize otherwise stable BHs.
+We conclude that, at least under static spherical symmetry, BH stability does not exhibit robustness.
+Two nearly identical BHs may exhibit completely different stability properties at late time.
+It will also show that horizon-induced redshift transforms near-horizon quantum fluctuations into classical-scale stochastic local deformations on the metric, which can trigger instability of the BH and cause it gradually to lose energy over time.
 In this paper, we use natural units $\hbar=c=G=1$.
-II Effective potential and its deformations
+II Effective potential and its alterations
 We take the original spacetime to be Schwarzschild BH, of which the line element is
-$\text{d}s^{2}=-h(r)\text{d}t^{2}+\frac{\text{d}r^{2}}{h(r)}+r^{2}\text{d}\Omega^{2},~{}~{}~{}h(r)=1-r_{h}/r\,,$ (1)
+$\text{d}s^{2}=-h\text{d}t^{2}+\frac{\text{d}r^{2}}{h}+r^{2}\text{d}\Omega^{2}\,$, (1)
+$h=1-\frac{r_{h}}{r}\,$, (2)
 where $r_{h}$ is the radius of event horizon, $\text{d}\Omega^{2}$ stands for line element of unit sphere.
-In time domain, we consider equation of motion with respect to the s-wave scalar perturbation $\psi(t,r_{*})$ in the Schwarzschild spacetime
+In time domain, we then consider equation of motion with respect to the s-wave scalar perturbation $\Psi=\psi(t,r_{*})/r$ as a test fieldin the Schwarzschild spacetime
 $\frac{\partial^{2}\psi}{\partial t^{2}}-\frac{\partial^{2}\psi}{\partial r_{*}^{2}}+V_{\mathrm{eff}}\psi=0\,.$ (3)
@@
 The $V_{\mathrm{RW}}$ is nonnegative, so the Schwarzschild BH is stable against the scalar perturbation, which has been confirmed inRegge and Wheeler (1957).
-We now add a deformation term into the effective potential to capture possible environmental effects presenting around the BH system.
+We now add an alteration term into the effective potential to capture possible environmental effects presenting around the BH system.
 Detailed discussion on how the environments change Eq.
-(2) will much involved and we here simply assume that they are all encoded into a deformation on effective potential
+(3) would be much involved and we here simply assume that they are all encoded into an alteration on effective potential
 $V_{\mathrm{eff}}=V_{\mathrm{RW}}(r_{*})+\epsilon V_{p}(r_{*}-a,t)\,$, (4)
@@
 (2024).
-In following discussions, we will specify $r_{h}=1$. We present our numerical results for solving Eq. (2) in time domain with initial data corresponding to Gaussian wave packet
-$\psi(0,r_{*})=\exp\left(\frac{r_{*}^{2}}{20}\right)\,,\quad\frac{\partial\psi}{\partial t}(0,r_{*})=0\,.$ (4)
-For spatial grid $r_{*}$, we adopt forth-order central difference method for spatial derivative and integrate the Eq.
+In the following discussions, we will specify $r_{h}=1$. We present our numerical results for solving Eq. (3) in time domain with initial data corresponding to Gaussian wave packet
+$\psi(r_{*},0)=\exp\left(-\frac{r_{*}^{2}}{20}\right)\,,\quad\frac{\partial\psi}{\partial t}(r_{*},0)=0\,.$ (5)
+For spatial grid $r_{*}$, we adopt fourth-order central difference method for spatial derivative and integrate the Eq.
 (3) using forth-order Rugge-Kutta method for $t$.
+At the horizon ($r_{*}\rightarrow$-$\infty$) and the infinity ($r_{*}\rightarrow\infty$) we impose the no reflect boundary conditions.
 Here we will consider three different types of potential $V_{p}(r_{*}-a,t)$.
 III Static negative bump
-The first one is the static “bump” potential $V_{p}(x,t)=V_{p}(x)$, where $V_{p}(x)$ has certain sign and compact support.
-When $V_{p}(x)$ is positive, it has been well known that such deformation will lead to “spectrum instability”.
+The first one is the “bump” static potential $V_{p}(x,t)=V_{p}(x)$, where $V_{p}(x)$ has certain sign and compact support.
+When $V_{p}(x)$ is positive, it has been well known that such alteration will lead to “spectrum instability”.
 However, this kind of “instability” is not the instability of BH itself, i.e., the system will still settle down into a static BH without scalar hair.
-Since the bumps are added by hand, there is not proper justification for a positive or negative sign.
+Since the bumps are added by hand, there is no proper justification for a positive or negative sign.
 Particularly, near the event horizon, the quantum effects of gravity and matter may lead to exotic matter and may raise negative “bump”.
-The negative “bump” can even appear if we impose dominate energy condition (see one example in appendix).
-The negative "bump" could resemble a tachyonic field Cardoso et al.
-(2009); Baibhav et al.
-(2023) or be inherent in various modified gravity theories, such as the Einstein-Gauss-Bonnet theory Cao et al.
-(2024).
-If the bump is close to a stable BH, the situation become complicated.
-We expect a competition with the absorption at the horizon.
-For strong enough negative bump, this feature might give rise to spacetime instabilitiesCardoso et al.
+The negative “bump” can even appear if we impose dominant energy condition (see one example in appendix).
+The negative "bump" could resemble a tachyonic field or be inherent in various modified gravity theories, such as the Einstein-Gauss-Bonnet theory.
+This feature might give rise to spacetime instabilities Cardoso et al.
 (2009); Baibhav et al.
@@
 (2024).
+If the bump is close to a stable BH, the situation becomes complicated.
+We expect a competition with the absorption at the horizon.
 Naively, one may expect that the BH will be still stable if the original BH is strongly enough but the negative “bump” is small enough.
-Surprisingly, our numerical results show that, if the infinitesimal negative “bump” is close to event horizon enough, it can turnover arbitrarily stable BH.
+Surprisingly, our numerical results show that, if the infinitesimal negative “bump” is close to event horizon enough, it can overturn an arbitrarily stable BH.
 We specify the negative bump to be Gaussian form around center $r_{*}=a$
 $V_{p}(r_{*}-a)=-\exp\left[-\frac{(r_{*}-a)^{2}}{16}\right]\,.$ (6)
-We numerically investigate how this negative bump change the long-time behavior of scalar perturbation when it is placed close to even horizon.
-Figure 1: Left panel: Time evolution of $\psi$ at $x=a$ for different values of $\epsilon$ and $a$.
-Here $t_{0}$ is defined as the first time when $\psi(a,t)$ is at its local maximum value.
+We numerically investigate how this negative bump changes the long-time behavior of scalar perturbation when it is placed close to even horizon.
+Figure 1: Left panel: Time evolutions of $\psi$ at $x=a$ for different values of $\epsilon$ and $a$.
+Here $t_{0}$ is defined as the first time at which $\psi(a,t)$ is at its local maximum value.
 Right panel: Relationship between $\epsilon$ and $|a_{c}|$.
@@
 shows the results as we move perturbed potential towards event horizon by choosing $\epsilon=10^{-2}$ and $4\times 10^{-3}$.
-It shows that for small value of $a$ where the deformations are confined near the photon ring, $\psi$ will decay to approach zero in the last stage, confirming that the BH is stable.
-However, for the case $\epsilon=10^{-2}$ we found that there exist a critical distance $a_{c}\approx-15$ of deformation to trigger the BH instability.
+It shows that for small value of $a$ where the alteration are confined near the photon ring, $\psi$ will decay to approach zero in the last stage, confirming that the Schwarzschild BH is stable.
+However, for the cases $\epsilon=10^{-2}$ we found that there exist a critical distance $a_{c}\approx-15$ of alteration to trigger the BH instability.
 When $a<a_{c}$, the perturbation field $\psi$ will experience exponential growth at late time.
-This exponential amplification signals an onset of instability, indicating that the perturbation field will drive the BH towards an unstable state during the later stage of its evolution.
+This exponential amplification signals an onset of instability, indicating that the perturbation field will drive the Schwarzschild BH towards an unstable state during the later stage of its evolution.
 If we decrease the amplitude $\epsilon$, then we find that the critical distance will also move towards horizon.
@@
 Our numerical results imply that the instability will always be triggered if the negative bump is sufficiently close to event horizon no matter how small the amplitude is.
+The “negative bump” we discussed here could destabilize the flat spacetime independently.
+However, the existence of BH will absorb energy and stabilize the spacetime.
+Thus, roughly speaking, the instability can be understood as the competition between the destabilizing caused by perturbed potential $V_{p}$ and the stabilizing caused by BH.
+One may expect that as long as the “negative bump” is sufficiently small and sufficiently close to the BH, the stable BH should dominate the stability of spacetime configuration, leading that entire spacetime configuration should tend to be stable.
+However, our results present a counterintuitive conclusion: no matter how small the bump is, it can overturn the stability of the entire spacetime, turning an originally stable Schwarzschild BH into an unstable one.
+Particularly, in the appendix B, we consider the situation that the background is not a BH but a star, then we find that infinitesimal negative bump cannot trigger the instability if it is placed sufficiently close to the star.
+Thus, the instability discovered here not only is caused by negative bump but also involves intrinsic property of the black hole geometry.
 IV stochastic local potential
@@
 More importantly, this kind of modification seemingly is inevitably even for an isolated BH without any matter surrounding it, since it can come from the local quantum fluctuations.
-Based on equivalent principle, the wavelength $\lambda_{p}$ and period $t_{p}$ of local quantum fluctuation measured by local inertial observers are Planck scales $\lambda_{p}=t_{p}\sim\mathcal{O}(l_{\rm P})$.
-These quantum fluctuations will lead that the geometry of spacetime varies in the local proper time scale $t_{p}$ and local proper distance scale $\lambda_{p}$.
-Usually, such extreme ultraviolet variaions of metric have no dramatic effect on classical physics.
-But the appearance of horizon makes thing different due to the strong redshift.
+Based on the equivalent principle, the wavelength $\lambda_{p}$ and period $t_{p}$ of local quantum fluctuation measured by local inertial observers are Planck scale $\lambda_{p}=t_{p}\sim\mathcal{O}(l_{\rm P})$.
+These quantum fluctuations will lead that the geometry of spacetime vary in the local proper time scale $t_{p}$ and local proper distance scale $\lambda_{p}$.
+Usually, such extreme ultraviolet variations of metric have no dramatic effect on classical physics.
+But the appearance of horizon makes things different due to the strong redshift.
 To illustrate that, we consider the quantum fluctuation happening in the near-event-horizon region $r=r_{h}+\delta r$ (measured in coordinate gauge (1)).
@@
 This means $\delta r/\sqrt{h(r)}\approx\delta r/\sqrt{h^{\prime}(r_{h})\delta r}\sim\mathcal{O}(\lambda_{p})$.
-One thus find that $T=\lambda\sim 1/h^{\prime}(r_{h})\sim\mathcal{O}(r_{h})$.
-Therefore, from the perspective of tortoise coordinate, near horizon local quantum fluctuations would induce metric or Regge-Wheeler potential to varies in infrared scale.
-It needs to emphasize again that, in the case of absent horizon or in the region far way from horizon, the local quantum fluctuations will have no essential contributions on classical physics in infrared scale.
+One thus finds that $T=\lambda\sim 1/h^{\prime}(r_{h})\sim\mathcal{O}(r_{h})$.
+Therefore, from the perspective of tortoise coordinate, near horizon local quantum fluctuations would induce metric or Regge-Wheeler potential to vary in infrared scale.
+It needs to emphasize again that, in the case of absent horizon or in the region far way from the horizon, the local quantum fluctuations will have no essential contributions on classical physics in infrared scale.
 In order to characterize the randomness of variation on Regge-Wheeler potential, we specify
 $V_{p}(x,t)=W(x)U(t)\,$, (8)
-where $W(x)$ is random function localizing around $x=0$ and $U(t)$ is a temporal random function. The function $W$ and $U$ satisfy
+where $W(x)$ is random function localizing around $x=0$ and $U(t)$ is a temporal random function. The function $W$ and $U$ are dimensionless functions and satisfy
 $\overline{W}=\int^{\infty}_{-\infty}W(x)\text{d}x=0,~\overline{U}=\lim_{t_{m}\rightarrow\infty}\frac{1}{t_{m}}\int^{t_{m}}_{-t_{m}}U(t)\text{d}t=0\,.$ (9)
@@
 In addition, the spectrum (after of Fourier transformations) of $W$ and $U$ should mainly distribute in the region that are not larger than the order $\mathcal{O}(1/r_{h})$.
-Figure 2: Left panel: Time evolution of $\psi$ at $x=a$ for different values of $\epsilon$ and $a$.
+Figure 2: Left panel: Time evolutions of $\psi$ at $x=a$ for different values of $\epsilon$ and $a$.
 Right panel: Relationship between $\epsilon$ and $|a_{c}|$.
@@
 V Low-frequency instability
-In above two examples, perturbed potentials $V_{p}$ themselves will trigger the instability if we turn off the Regge-Wheeler potential $V_{\text{RW}}$.
-Thus, roughly speaking, the instability can be understood as the competition between the destabilizing caused by perturbed potential $V_{p}$ and the stabilizing caused by BH.
-The following example will show that this is not the whole story.
+As we previously discussed, the stochastic local potential not only comes from the quantum fluctuations but also classical fluctuations.
+In Sec.
+IV, we focus on the quantum fluctuations with frequency near Planck scale, which will become $\mathcal{O}(1/r_{h})$ due to the extreme redshift near BH horizon.
+Since the proper frequency of classical fluctuations are much lower than the quantum fluctuations, we thus consider the low frequency $w\ll{\cal O}(1/r_{h})$ limit for the case of classical fluctuations near the BH horizons.
 We specify the perturbed potential $V_{p}$ to be
 $V_{p}(r_{*}-a,t)=\exp\left[-\frac{(r_{*}-a)^{2}}{\sigma^{2}}\right][\sin(wt)+\sin(\pi wt)]\,$, (13)
-i.e, we consider a time-dependent perturbed potential that has Gaussian distribution in space but pseudo-random in time. We here focus on the situation $w\ll\mathcal{O}(1/r_{h})$.
-Figure 3: Left panel: Time evolution of $\psi$ at $x=a$ in different situations with $w=0.02$ and $\sigma=4$.
+i.e, we consider a time-dependent perturbed potential that has Gaussian distribution in space but pseudo-random in time.
+Here we set $w\ll 1$.
+In this situation, we will show that deformation potential itself will stabilize the flat background but can trigger the instability if it is closed to a BH.
+Figure 3: Left panel: Time evolutions of $\psi$ at $x=a$ in different situations with $w=0.02$ and $\sigma=4$.
 Right panel: Relationship between $\epsilon$ and $|a_{c}|$ for $w=0.001$ and $\sigma=4$.
@@
 for sufficiently small $\epsilon$.
-Our numerical results suggest that, for the case $w\rightarrow 0^{+}$ the infinitesimal $\epsilon$ can still trigger the instability if the perturbed potential is sufficiently closed to the horizon.
-This effect appears unique to black holes as these same potentials do not destabilize Minkowski spacetime.
+Our numerical results suggest that, for the case $w\rightarrow 0^{+}$ the infinitesimal $\epsilon$ can still trigger the instability if the perturbed potential is sufficiently close to the horizon.
+This effect appears unique to BHs as these same potentials do not destabilize Minkowski spacetime.
 VI conclusion
@@
 These classical and quantum perturbations induce stochastic deformations in spacetime geometry.
-This study addresses a critical question: whether nominally stable BHs retain their stability when the geometry are deformed by such small perturbations.
+This study addresses a critical question: whether nominally stable BHs retain their stability when the geometry is deformed by such small perturbations.
 To model this effect, we map the deformations onto modifications of the effective potential governing s-wave scalar fields.
@@
 Our findings suggest that the stability of a stable BH is not robust against the tiny deformation of near-horizon geometry.
-Infinitesimal deformation of near-horizon geometry can turnover the stability of a stable BH.
+Infinitesimal deformation of near-horizon geometry can overturn the stability of a stable BH.
 Due to the inevitable fluctuations of near horizon geometry, this intrinsic instability fundamentally constrains our ability to predict BH dynamics over extended timescales.
 As a toy model and very preliminary research, we here only considered static spherically symmetric case and numerically simulated the system in time domain.
-It would be interesting to consider Kerr black hole and study how spin would affect the phenomenon discovered here.
-Particularly, is there any relationship to turbulence of black holes discovered in Ref.
+It would be interesting to consider Kerr BH and study how spin would affect the phenomenon discovered here.
+Particularly, is there any relationship to turbulence of BHs discovered in Ref.
 Yang et al.
@@
 [^1] The analysis of frequency domain for static negative bump and static stochastic potential will be addressed in an upcoming publication.
-Why does bump potential modulated by a time function in low frequency limit stabilize the flat spacetime but destabilize the black hole?
 More realistically, could we quantitatively estimate the disturbance intensity $\epsilon$ generated by the actual environment (such as accretion disks, dark matter halos) or quantum fluctuations, and whether the position $a$ satisfies the condition of $a<a_{c}$?
-We hope these questions could be addressed in the future through numerical simulations and further theoretical works.
+We hope these questions could be addressed in the future through numerical simulations and further theoretical work.
 At the classical level, we emphasize that our results do not imply instability of the vacuum Schwarzschild solution, as no source exists to generate stochastic deformations on metric.
-For a BH surrounded by matter, our results implies that the system cannot keep static configuration for a long time in astrophysical settings, since uncontrollable stochastic matter fluctuations near the horizon will lead stochastic deformations on the geometry and inevitably trigger the instability identified here.
+For a BH surrounded by matter, our results imply that the system cannot keep static configuration for a long time in astrophysical settings, since uncontrollable stochastic matter fluctuations near the horizon will lead to stochastic deformations on the geometry and inevitably trigger the instability identified here.
 The timescale that this instability becomes considerable depends on amplitude of the deformation.
-The probe scalar field studied here can act as a conduit, transferring energy from the environment to the BH and spatial infinity.
+The probe scalar field studied here can act as a conduit, transferring energy from the environment to the BH and spatial infinity (See also Appendix.
+C).
 Consequently, the final state evolves toward a Schwarzschild BH in vacuum, with increased horizon area and energy.
-However, if we posit an underlying quantum-gravitational framework beneath classical general relativity, our analysis indicates that even vacuum Schwarzschild BHs become unstable.
-This suggests that Schwarzschild BHs in a quantum-gravitational framework would gradually lose energy over time, potentially offering a new perspective on BH evaporation.
+However, if we posit an underlying quantum-gravitational framework beneath classical general relativity, it becomes very interesting.
+We found that the quantum fluctuations could destabilize the Schwarzschild BH, implying that the scalar perturbations will exponentially grow with time and keep transferring energy to infinity.
+Will our finding be a potential mechanism for a new perspective on BH evaporation?
+This is a very interesting question.
 Further investigation is needed to establish a direct connection and we hope some useful advances will be obtained in the future.
@@
 This work is also supported by the Guangxi Talent Program (“Highland of Innovation Talents”).
-Appendix A Negative bump potential and dominate energy condition
-In this appendix, we will give a concrete example that Regge-Wheeler potential has a local negative bump but dominate energy condition is satisfied.
+Appendix A Negative bump potential and dominant energy condition
+In this appendix, we will give a concrete example that Regge-Wheeler potential has a local negative bump but dominant energy condition is satisfied.
 Let us first begin from general static spherical asymptotically flat BH in tortoise coordinates, of which the metric reads
@@
 We here assume that the event horizon locates at $x=$-$\infty$, where we have $r(-\infty)=r_{h}$ and $h(-\infty)=0$.
-The function $h(x)$ has asymptotically expression $h(x)=1-2M/x+\cdots$ with the total mass $M$ of the spacetime and $r^{\prime}(\infty)=1$.
+The function $h(x)$ has asymptotical expression $h(x)=1-2M/x+\cdots$ with the total mass $M$ of the spacetime and $r^{\prime}(\infty)=1$.
 The Regge-Wheeler potential $V_{\mathrm{RW}}$ for s-wave scalar field then reads
@@
 In static spherically symmetric case, the energy momentum tensor ${T^{\mu}}_{\nu}$ has a form ${T^{\mu}}_{\nu}=\text{diag}[-\rho(x),p_{x}(x),p_{T}(x),p_{T}(x)]$.
-The dominate energy condition then requires
+The dominant energy condition then requires
 $\rho\geq|p_{x}|,~\rho\geq|p_{T}|\,.$ (17)
@@
 The values of $\rho$ and $p_{x}$ could be considerably large near horizon.
-Particularly, if an astrophysical BH is surrounded by sufficient matter localization around $x\sim a\ll-r_{h}$ and the dominate energy condition is satisfied, then it is possible
+Particularly, if an astrophysical BH is surrounded by sufficient matter localization around $x\sim a\ll-r_{h}$ and the dominant energy condition is satisfied, then it is possible
 $4\pi r^{2}(\rho-p_{x})>1-r^{\prime 2}/h\,$ (21)
@@
 as an example.
-To solve the Einstein equation, we note that the system is in vacuum and the solution is a Schwarzschild BH with Schwarzschild radius $r_{h}$ in the region $x<a-1/8$, of which the metric reads
-$x=r_{h}+r_{h}\ln(1-r_{h}/r),~{}~{}r^{\prime}=h=1-r_{h}/r\,.$ (22)
-We then integrate equations (17) and (18) into infinity with the initial values at $x=a-1/8$ given according to Eq. (22).
+To solve the Einstein equation, we note that the system is in vacuum and the solution is a Schwarzschild BH with Schwarzschild radius $r_{h}$ in the region x<a1/8x<a-1/8, of which the metric reads
+$x=r+r_{h}\ln(1-r_{h}/r),~~r^{\prime}=h=1-r_{h}/r\,.$ (23)
+We then integrate equations (18) and (19) into infinity with the initial values at x=a1/8x=a-1/8 given according to Eq. (23).
 Figure 4: Regge-Wheeler potential for $a=-20,r_{h}=1$ and $b=0.2$. Here $x_{m}\approx-11.09$ is the position where $V_{\mathrm{RW}}$ reaches its maximum.
@@
 We see that there is negative bump around $x=a$.
-Thus, the negative bump potential can appear even if dominate energy condition is satisfied.
+Thus, the negative bump potential can appear even if dominant energy condition is satisfied.
 We expect that, in general, localized spherical matter shells will introduce similar bumps in the Regge-Wheeler potential, possibly causing instabilities discussed in our main text.
+Appendix B Infinitesimal negative bump in static spherical star
+In this appendix, we consider a static ideal star as background, then we find that infinitesimal negative bump cannot trigger the instability if it is placed sufficiently close to the star.
+Since an astrophysical star involves much complicated details of its interior, we here just us a toy model and only regard the star as a horizonless geometry with nonzero total mass.
+We consider a regular Hayward star.
+The metric of the Hayward star is  Hayward (2006)
+$\text{d}s^{2}=-h(r)\text{d}t^{2}+\frac{\text{d}r^{2}}{h(r)}+r^{2}\text{d}\Omega^{2}\,,\quad h=1-\frac{2Mr^{2}}{r^{3}+2b^{2}M}\,.$ (24)
+When $M<3\sqrt{3}/4b$, there is no horizon and the metric thus describe a regular star.
+For without the loss of generality, we set that $M=1,b=2$ as a concrete example.
+One important difference is that the tortoise coordinate $r_{*}$ here is defined as
+$r_{*}=\int_{0}^{r}\frac{\text{d}x}{h(x)}\in[0,\infty)$ (25)
+This leads fundamental changes: For BH, “infinitely close to horizon” means $r_{*}\rightarrow$-$\infty$; For a star, “infinitely close to star” means $r_{*}$ is still finite and can at most decrease to zero.
+This difference is universal for general spherically symmetric stars and BHs.
+Figure 5: Time evolutions of $\psi$ at $x=a$ in different situations with $\epsilon=0.01$ and 0.001
+In Fig.
+5, we numerically shows two cases of $a=0$: $\epsilon=0.01$ and $0.001$.
+We choose $a=0$ since this stands for the case that is closest to the star.
+Clearly, sufficiently small negative bump cannot trigger the instability when it is “sufficiently close to the star”.
+[^2] When the negative bump is sufficiently far way from the star, the instability can still be triggered but this is the effect of negative bump in flat space.
+Appendix C Energy flow of the scalar field
+To explain how the probe scalar field would serve as the conduit, we can compute the energy fluxes towards event horizon and infinity. For a light massless scalar field, the energy momentum tensor
+$T^{\mu\nu}=\nabla^{\mu}\Psi\nabla^{\nu}\Psi-\frac{1}{2}g^{\mu\nu}(\nabla\Psi)^{2}\,$, (26)
+where $g^{\mu\nu}$ denotes the metric of Schwarzschild black hole background and $\Psi=\psi/r$.
+It needs to note that asymptotical flatness requires $\Psi$ to be zero at infinity but $\psi$ could be finite.
+Since the background spacetime is assumed to be static with timelike Killing vector $\xi^{\mu}=(\partial/\partial t)^{\mu}$, then we can define the an energy current $J^{\mu}=-T^{\mu\nu}\xi_{\nu}$ and it satisfies $\nabla_{\mu}J^{\mu}=0$.
+In order to study the energy fluxes at horizon and infinity, let us consider a family of spacelike hypersurfaces $\{\Sigma_{s}\}$ outside event horizon, which gives a foliation on the spacetime region between $r_{h}$ and $r=r_{0}\gg r_{h}$.
+We assume that this foliation preserve the spherical symmetry.
+The boundaries of every $\Sigma_{s}$ are given by $r=r_{h}$ and $r=r_{0}$.
+The energy in one hypersurface $\Sigma_{s}$ is given by
+$E_{s}=\int_{\Sigma_{s}}J^{\mu}\text{d}\Sigma_{\mu}\,.$ (27)
+Here $\text{d}\Sigma_{\mu}$ is the future-toward directed volume element. Let us consider two different hypersurface $\Sigma_{1}$ and $\Sigma_{2}$, see Fig. 6.
+Figure 6: Penrose diagram on the exterior of horizon and a family of spacelike hypersurfaces $\{\Sigma_{s}\}$.
+Here $\mathcal{N}$ is boundary belong to horizon and $S$ is the boundary at $r=r_{0}\gg r_{h}$.
+In the limit $r_{0}\rightarrow\infty$, we have $S\subset i_{0}$.
+It should be note that $\xi^{\mu}$ is ‘outward directed’ normal to $\mathcal{N}$, as determined by continuity.
+The difference of energies between these two hypersurfaces is
+$\begin{split}&E_{2}-E_{1}=\int_{\Sigma_{2}}J^{\mu}\text{d}\Sigma_{\mu}-\int_{\Sigma_{1}}J^{\mu}\text{d}\Sigma_{\mu}\\ &=\oint_{\partial V}J^{\mu}\text{d}\Sigma_{\mu}-\left(\int_{\mathcal{N}}J^{\mu}\text{d}\Sigma_{\mu}+\int_{S}J^{\mu}\text{d}\Sigma_{\mu}\right)\,.\end{split}$ (28)
+Here $\mathcal{N}$ is boundary belong to horizon, $S$ is the boundary at $r=r_{0}$ and $V$ is the spacetime region surrounded by $\Sigma_{1},\Sigma_{2},\mathcal{N}$ and $S$.
+The hypersurface $S$ is define by $r=r_{0}$ and $t\in[t_{1},t_{2}]$.
+The null hypersurface $\mathcal{N}$ is defined by $r=r_{h}$ and $v\in[v_{1},v_{2}]$, where $(\partial/\partial v)^{\mu}$ is the future-directed generator of event horizon.
+At event horizon, $\xi^{\mu}$ is both normal and tangent to event horizon and we take $(\partial/\partial v)^{\mu}=\xi^{\mu}$.
+After using Gaussian theorem and $\nabla_{\mu}J^{\mu}=0$, we then have
+$E_{2}-E_{1}=-\left(\int_{\mathcal{N}}J^{\mu}\text{d}\Sigma_{\mu}+\int_{S}J^{\mu}\text{d}\Sigma_{\mu}\right)\,.$ (29)
+The above two integrations describe the energy fluxes at horizon and $r=r_{0}$, respectively.
+Since $r=r_{0}$ is a timelike surface and its unit outer normal vector reads $n^{\mu}=(\partial/\partial r)^{\mu}$ (we assume that $r_{0}\gg r_{h}$).
+The energy flux at $r=r_{0}$ then is given by
+$\begin{split}\int_{S}J^{\mu}\text{d}\Sigma_{\mu}&=-4\pi r_{0}^{2}\int_{t_{1}}^{t_{2}}\text{d}t\dot{\Psi}\partial_{r}\Psi|_{r=r_{0}}\\ &=\left.-4\pi\int_{t_{1}}^{t_{2}}\text{d}t\dot{\psi}\left(\frac{\partial\psi}{\partial r}-r^{-1}\psi\right)\right|_{r=r_{0}}\,.\end{split}$ (30)
+The “no-reflect boundary condition” at the infinity shows that
+$\left.\dot{\psi}+\frac{\partial\psi}{\partial r}\right|_{r\rightarrow\infty}=0\,$,
+and asymptotical flatness requires $\psi/r=\Psi\rightarrow 0$.
+The energy flux at $S$ then is positive and is toward infinity.
+Since $\mathcal{N}$ is a null surface and its outer normal vector reads $\xi^{\mu}$ (we assume that $r_{0}\gg r_{h}$).
+The energy flux at $\mathcal{N}$ then is given by
+$\begin{split}&\int_{\mathcal{N}}J^{\mu}\text{d}\Sigma_{\mu}=-\int_{\mathcal{N}}J^{\mu}\xi_{\mu}\text{d}\Sigma\\ &=4\pi r_{h}^{2}\int_{v_{1}}^{v_{2}}\text{d}t\dot{\Psi}\partial_{t}\Psi|_{r=r_{0}}=4\pi\int_{v_{1}}^{v_{2}}\text{d}v\dot{\psi}^{2}\,.\end{split}$ (31)
+Here we have used the fact that $\xi^{\mu}$ is null vector at horizon. The energy flux towards interior of event horizon then is also positive.
+In the stable case, any fluctuation of test scalar field will decay to zero fast, so energy fluxes at $\mathcal{N}$ and $S$ will also decay to zero fast.
+However, in the situations discussed in main text, the fluctuation of test scalar field will exponentially increase, so energy fluxes at $\mathcal{N}$ and $S$ will be more and more.
+Therefore the scalar field acts as a conduit, transferring energy from the environment into the black hole and out to infinity.